OpenStudy (thatonegirl_):

If the graph of the function defined on [-3,3] by f(x)=x^2+ax+b has an absolute minimum at (-1,-3), determine the value of f(1).

OpenStudy (irishboy123):

complete the square for a quadratic....? or calculus? what you learning? :-)

OpenStudy (thatonegirl_):

calculus :)

OpenStudy (thatonegirl_):

i can't remember how to do tht part though lol

OpenStudy (irishboy123):

lol!!

OpenStudy (irishboy123):

small detail

OpenStudy (thatonegirl_):

Lol yeah

OpenStudy (irishboy123):

so we can compute \(f(1)\) like now but maybe if we figure out what \(f'\) is, it might also add info

OpenStudy (thatonegirl_):

So i would take the derivative of that equation?

OpenStudy (irishboy123):

totally!

OpenStudy (thatonegirl_):

sweet so its 2x+a

OpenStudy (irishboy123):

and it's a min at (-1, -3) apparently

OpenStudy (thatonegirl_):

yup

OpenStudy (irishboy123):

OK!! lol so you've got that it's a min, which means something about the 2nd deriv, i reckon

OpenStudy (thatonegirl_):

Isn't 2nd deriv for concavity?

OpenStudy (irishboy123):

i was meandering it goes through (3,3) so you have your other equation too

OpenStudy (thatonegirl_):

wait how do you know it goes through that point?

OpenStudy (eliesaab):

Solve f'[-1]=0 and f[-1]=3 to find a and b replace and find f[1]

OpenStudy (eliesaab):

Sorry f[-1]=-3

OpenStudy (irishboy123):

\( f=x^2+ax+b \) \(f(1) = 1 + a + b\) \( f'=2x+a\) \( f'(-1)=2(-1)+a = 0 \implies a = 2\) \(\implies f=x^2+2x+b \) \(\implies 3 =(-3)^2+2(-3)+b \) \(\implies b = 0 \) \(\implies f=x^2+2x+b \)

OpenStudy (eliesaab):

Who said that f(3)=3?

OpenStudy (eliesaab):

All what we know is f(-1)=-3

OpenStudy (thatonegirl_):

oh yes

OpenStudy (irishboy123):

yes?

OpenStudy (eliesaab):

If you do what I said above you find \[ f(x)=x^2+2 x-2\\ f(1)=1 \]

OpenStudy (thatonegirl_):

and a would equal 2. \[f'(x)=2x+a\] \[f'(-1)=-2+a\] \[-2+a=0\] \[a=2\]

OpenStudy (thatonegirl_):

how did u find b again? sorry

OpenStudy (irishboy123):

@eliesaab is not happy yet better at math than I i reckon we roll that way :-)

OpenStudy (thatonegirl_):

Lol no I appreciate both of your help.

OpenStudy (thatonegirl_):

nvm i got b

OpenStudy (thatonegirl_):

thank you guys!!

OpenStudy (irishboy123):

@eliesaab we also are told the slope :-)

OpenStudy (irishboy123):

as in the min at (-1, -3)

OpenStudy (thatonegirl_):

I got the answer, thanks! :D

OpenStudy (irishboy123):

yes @Thatonegirl_ you don't have to worry!! good stuff close it and forget about it